# Advancing Brainwave-Based Biometrics: A Large-Scale, Multi-Session Evaluation

### v1

@@
Abstract.
The field of brainwave-based biometrics has gained attention for its potential to revolutionize user authentication through hands-free interaction, resistance to shoulder surfing, continuous authentication, and revocability. However, current research often relies on single-session or limited-session datasets with fewer than 55 subjects, raising concerns about generalizability and robustness. To address this gap, we conducted a large-scale study using a public brainwave dataset of 345 subjects and over 6,000 sessions (averaging 17 per subject) recorded over five years with three headsets. Our results reveal that deep learning approaches outperform classic feature extraction methods by 16.4% in Equal Error Rates (EER) and comparing features using a simple cosine distance metric outperforms binary classifiers, which require extra negative samples for training. We also observe EER degrades over time (e.g., 7.7% after 1 day to 19.69% after a year). Therefore, it is necessary to reinforce the enrollment set after successful login attempts. Moreover, we demonstrate that fewer brainwave measurement sensors can be used, with an acceptable increase in EER, which is necessary for transitioning from medical-grade to affordable consumer-grade devices. Finally, we compared our findings with prior work on brainwave authentication and industrial biometric standards. While our performance is comparable or superior to prior work through the use of Supervised Contrastive Learning, standards remain unmet. However, we project that achieving industrial standards will be possible by training the feature extractor with at least 1,500 subjects. Moreover, we open-sourced our analysis code to promote further research.
@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Biometricsccs: Security and privacy Usability in security and privacy
1. Introduction
@@
Authentication is the cornerstone of securing digital interactions. While passwords have long been the dominant method, biometrics are increasingly preferred. According to the 2024 FIDO Survey [^1] https://fidoalliance.org/wp-content/uploads/2024/10/Barometer-Report-2024-Oct-29.pdf of 10,000 consumers, 28% preferred biometrics for signing into online accounts, apps, and smart devices, making it the most popular choice. Beyond established methods—such as fingerprint and face recognition— emerging biometric technologies are being explored to enhance usability and security. Brainwave authentication has gained attention for its hands-free operation, which enhances usability in devices like XR systems (Lin et al., 2018; Fallahi et al., 2024a). It resists shoulder surfing by relying on distinct neural activity patterns rather than observable actions or physical characteristics (Arias-Cabarcos et al., 2023). Additionally, brainwave-based systems can enable continuous authentication, improving security by monitoring user identity in active tracking (Nakanishi and Yoshikawa, 2015). Furthermore, research indicates that in cases of brainwave sample leakage, changing the stimulus can effectively reset the authentication pattern, thereby maintaining security (Lin et al., 2018). However, brainwave authentication is still in its early stages, and both permanence and performance remain significant challenges.
@@
Prior studies on brainwave authentication mainly rely on limited public datasets or small, privately collected datasets. Consequently, studies either focus on single-session datasets (Schons et al., 2018; Bidgoly et al., 2022; Arias-Cabarcos et al., 2023; Fallahi et al., 2023) or small-scale multi-session datasets (Das et al., 2016; Wu et al., 2018; Maiorana, 2021; Debie et al., 2021). While single-session studies completely neglect session and temporal effects (Chaurasia et al., 2024; Huang et al., 2022), multi-session studies are limited to at most 55 (Debie et al., 2021) subjects and 270  (Maiorana and Campisi, 2017) sessions (6 per subject), increasing the risk of overfitting and restricting the ability to generalize results and identify key factors influencing performance and reliability over time. Moreover, the lack of publicly accessible code and data hinders reproducibility and further development, in contrast to the face recognition field, which benefits from open data and collaborative research.
@@
To address this gap and better understand how different parameters affect the performance of brainwave authentication over time, we used the recently published PEERS dataset (Kahana et al., 2024), which comprises data from 345 subjects and over 6,000 sessions spanning five years. This dataset contains more than six times the number of subjects and 22 times the number of sessions compared to previous brainwave authentication evaluations. Our main contributions are as follows:
  • Benchmark (Sec 4.1): To identify the best-fit pipeline for multi-session authentication, we performed a comprehensive investigation of both classical and deep learning-based feature extraction approaches. Our results showed that pairwise-based learning methods, delivered superior performance for brainwave-based user authentication, reducing the EER from 27.3% to 10.9%. We also examined distance-based and binary classification-based feature comparison methods for brainwave data. While both approaches exhibited comparable performance, we chose the distance-based method because it does not require negative samples, making it more practical.
@@
  • Analysis (Sec 4.2): To better understand the effective parameters influencing performance, we investigated the effects of test size, session intervals, and hardware type, as well. Our analysis reveals several key insights. First, a larger number of subjects in the test set does not necessarily increase the authentication error. However, smaller evaluation datasets introduce higher uncertainty due to the limited number of observations. Second, examining the time intervals between sessions shows that authentication errors increase over time: approximately 1% after two weeks, 4% after three months, and 12% after twelve months of enrollment. This indicates that, depending on the use case, it was beneficial to update the enrollment dataset after successful logins to reduce long-term error rates. Third, visualization of the embedding space reveals that samples from the same session are relatively close, which highlights the effect of sessions on the learning process, potentially caused by environmental noise, device setup, and changes in the subject over time. Finally, the analysis of the three devices used for dataset collection shows that cross-hardware authentication is only feasible when a sufficient number of subjects in the training set include cross-hardware sessions with the same devices; otherwise, the error would be too high.
@@
  • Enhancing Performance (Sec 4.3): We investigated the trade-off between verification time and performance, reducing the error rate from 10.78% to 6.87% at the cost of increased verification time, which ranged from 1 to 32 seconds. Additionally, we examined the trade-off between enrollment sessions and error rates, observing that after the second enrollment session, the improvement in error rate diminished.
@@
  • Channel Reduction (Sec 4.4): To evaluate the system’s performance under conditions resembling real-world scenarios, we reduced the number of channels in the PEERS dataset—collected using medical-grade devices—from 93 to 14, 7, and 4. These reductions correspond to sensor placements commonly found in consumer-grade EEG devices. The results showed a relatively small increase in error compared to the significant reduction in channels. For instance, the error increased from 10.78% EER with 93 channels to 13.97% with 14 channel. This modest increase can be attributed to the high correlation in EEG data, which helps mitigate the impact of channel reduction on performance.
  • Industrial Standard Comparison (Sec 4.5): We compare our results as the state of the art in brainwave authentication to industrial standards for biometric implementation, such as the National Institute of Standards and Technology (NIST) and the International Organization for Standardization (ISO). While our results outperform or are comparable to other brainwave authentication works, despite using a more challenging test set comprising 100 subjects, we observe that at the security levels expected in industrial standards, the FRR is high. However, our complementary experiments showed a linear relationship between the error rate and logarithm of the number of subjects in the feature extraction training set. Therefore, we predict that with at least 1,500 subjects, the standards could be met at a usable FRR.
@@
  • Open Source: We published our source code, which, along with the public dataset, makes results easily reproducible and facilitates future error rate reductions by other researchers  [^2] https://github.com/kit-ps/NeuroShield/.
2. Background and Related Work
@@
EEG measures electrical activity produced by neuronal firings in the human brain, typically recorded at the scalp using specialized sensors (Niedermeyer and da Silva, 2005; Hu and Zhang, 2019). German physiologist and psychiatrist Hans Berger recorded the first human EEG in 1924 for medical purposes, which remains one of its primary applications (Gupta et al., 2022; Adeli and Ghosh-Dastidar, 2010; Haas, 2003). Beyond medical use, EEG technology has advanced to facilitate direct communication between the brain and external devices, known as Brain-Computer Interfaces (BCI) (Kawala-Sterniuk et al., 2021). Recently, EEG data has also been found to include unique features linked to user identity, making it a promising method for user authentication (Gui et al., 2019).
@@
Biometric authentication modalities should meet certain requirements to enable widespread usage. These requirements include universality (the modality must be present in all individuals), distinctiveness (it must uniquely identify an individual), permanence (it should remain consistent over time), and collectability (it must be measurable with available technology). Additional factors include performance (accuracy, speed, and resource efficiency), acceptability (user willingness to utilize the system), and circumvention resistance (difficulty in forging or bypassing) (Rui and Yan, 2018).
@@
EEG is universal because each person’s brain produces EEG signals (Marcel and Millán, 2007). These signals exhibit distinctiveness due to individual brainwave patterns, making them uniquely identifiable (Kaliraman et al., 2019). Advances in technology are rapidly improving collectability, with new EEG headsets entering the market annually, offering better signal quality and usability (Yang and Deravi, 2017). However, without real-time implementation, it is difficult to comprehensively assess acceptability and circumvention resistance, though some research attempts have been made in this direction (Fallahi et al., 2024b; Wang et al., 2022).
@@
Currently, the primary focus of research is on performance and permanence, as these are challenging aspects of EEG-based authentication. The two are closely interrelated, with performance metrics reflecting the system’s ability to maintain consistent accuracy over time. Factors such as hair growth, slight variations in sensor placement on the scalp, brain states (e.g., stress), and environmental conditions (e.g., noise) can significantly impact both performance and permanence in EEG-based authentication. However, most research studies evaluate the performance of brainwave authentication using single-session datasets (Schons et al., 2018; Bidgoly et al., 2022; Arias-Cabarcos et al., 2023; Fallahi et al., 2023), which neglect session invariance (permanence) completely. This limitation raises concerns that models may overfit to session-specific characteristics rather than capturing subject-unique features. Additionally, some studies (Debie et al., 2021; Chaurasia et al., 2024) consider short time intervals, such as within a few days, which may only partially account for the aforementioned issues. While a limited number of papers explore time intervals spanning more than a week, these studies (Das et al., 2016; Wu et al., 2018; Maiorana and Campisi, 2017; Maiorana, 2021) rely on non-public datasets and omit sharing their source code, hindering reproducibility and further research. Moreover, such studies have been conducted with a relatively small sample size and limited sessions (up to 54 subjects (Debie et al., 2021) and 270 sessions (Maiorana and Campisi, 2017)), which increases the risk of overfitting and restricts the generalizability of the results.
@@ 3. Method
Biometric systems, including brainwave authentication, involve four key stages: data collection, preprocessing, feature extraction, and feature comparison, and, like other biometric systems, they operate in two phases: enrollment and verification. In the enrollment phase, brainwave signals are collected from the user to generate a reference profile, known as a template. During verification, performed after enrollment, brainwave signals are captured again and compared with the reference profile to confirm the user’s identity. Due to the high-dimensional and noisy nature of biometric data, including brainwaves, direct comparison of raw samples is impractical. Preprocessing is therefore applied post-data collection to eliminate noise and extract relevant features that preserve subject-specific uniqueness while reducing dimensionality. A similarity metric is subsequently employed to compare these processed features and determine whether to accept or reject the authentication request.
To address the challenges posed by the noisy and high-dimensional nature of brainwave data, some systems incorporate external stimuli during data collection (Alzahab et al., 2022; Arias-Cabarcos et al., 2023; Fallahi et al., 2023). These stimuli, such as visual or auditory signals, elicit distinct brainwave responses, allowing the system to timestamp and isolate signal segments directly related to the response. This approach enhances the signal-to-noise ratio and improves the reliability of the feature extraction process, ultimately strengthening the accuracy of the authentication system.
3.1. Dataset
@@
Our study utilizes the Penn Electrophysiology of Encoding and Retrieval Study (PEERS) dataset (Kahana et al., 2024), provided by the University of Pennsylvania. Collected between 2010 and 2020, the dataset focuses on exploring EEG correlates of memory processes, particularly during tasks involving word memorization and recall. We used memorization tasks for user authentication, as the timestamps of word presentation available in the metadata can be considered as stimuli, while recall timestamps are not available. The dataset encompasses approximately 8.7 TB of EEG data, comprising recordings from 345 subjects across more than 6000 sessions. The participants range in age from 17 to 85 years, with an average age of 26.98 $\pm$ 15.86 years. The dataset includes 161 female and 136 male participants, while sex information for 48 subjects is unavailable. The PEERS dataset is particularly suited for our research as it includes multiple sessions per subject, enabling the investigation of inter-session variability, a critical factor for ensuring the robustness of real-world biometric systems. Data collection was conducted using three distinct medical grade headsets: the 129-channel Geodesic Sensor Net (GSN 200 model) [^2] https://med.stanford.edu/content/dam/sm/lucasmri/documents/16_0824_EGI_geodesic_sensor_net.pdf , the 129-channel HydroCel Geodesic Sensor Net. [^2] https://med.stanford.edu/content/dam/sm/lucasmri/documents/16_0824_EGI_geodesic_sensor_net.pdf, and the 128-channel BioSemi [^3] https://www.biosemi.com/headcap.htm headcap using the Biosemi ActiveTwo acquisition system. These diverse devices further enhance the dataset’s relevance by reflecting real-world variability in EEG acquisition systems.
3.2. Preprocessing
@@
As the channel layout of the different headsets varied, we unified them by mapping the channels to a well-known extended version [^4] https://mne.tools/stable/auto_tutorials/intro/40_sensor_locations.html of the ten-twenty electrode system (Silverman, 1963), which includes 93 channels. Subsequently, the EEG data channels were reordered for each recording to align with the expected order during evaluation across all three devices. Following this, and in line with best practices for preprocessing in brainwave authentication (Chaurasia et al., 2024; Arias-Cabarcos et al., 2023; Peng, 2019; Maiorana, 2021), a common average reference is applied to the data, which helps in reducing noise and improving signal quality (Maiorana, 2021). Then, a bandpass filter is applied to retain frequencies between 0.1 Hz and 60 Hz. This filtering step removes low-frequency drift and high-frequency noise, preserving the components of the signal most relevant for EEG analysis (Peng, 2019; Felja et al., 2023). Finally, we extracted 1-second samples based on stimulus timestamps. The sample data was saved along with the session date and hardware type. In the paper, we refer to the oldest session for each subject as the enrollment sample, while the remaining sessions are referred to as verification samples, unless otherwise specified.
3.3. Feature Extraction
After preprocessing, we have samples that include 46,500 values (93 * 500). This high-dimensional samples, considering the inherent variability of EEG data, cannot be compared directly. Therefore, we need to extract features related to the identity of the subject. There are two main approaches: the first involves classic feature extraction, which relies on predefined formulas derived from expert knowledge in the field. The second approach employs deep learning techniques for feature extraction, allowing models to automatically learn representations from the data.
We select Power Spectral Density (PSD) as the classic feature extraction technique because it effectively transfers signals from the time domain to the frequency domain, capturing unique patterns linked to individual brain dynamics. Consequently, PSD is a widely utilized feature extraction technique in brainwave-based authentication systems (Khalil et al., 2024; Ong et al., 2018; Thomas and Vinod, 2018; Arias-Cabarcos et al., 2023; Chaurasia et al., 2024). PSD applies Fourier transformation to quantify the distribution of signal power across different frequency components (Welch, 1967). The resulting frequency spectrum can then be analyzed either within predefined frequency bands—such as delta, theta, alpha, beta, and gamma—or across the full spectrum, depending on the specific application requirements (Arias-Cabarcos et al., 2023).
For deep learning feature extraction, we used a metric learning approach, which is a type of machine learning paradigm designed to learn a similarity or distance function that maps input data to a feature space where similar instances are closer together and dissimilar instances are farther apart (Kulis et al., 2013). This method is particularly suitable for biometric authentication as it ensures that features of the same identity are clustered closely while features from different identities are well-separated (Schroff et al., 2015; Wang et al., 2017). Metric learning approaches can be categorized into pairwise-based and softmax-based methods. Pairwise-based loss functions optimize embedding distances by focusing directly on pairs or triplets of samples, increasing the number of training instances derived from the dataset. This property makes them particularly effective for small datasets. In contrast, softmax-based methods optimize class separability by mapping embeddings into distinct clusters, typically requiring larger datasets for effective generalization (Horiguchi et al., 2019).

### v2

@@
Abstract
The field of brainwave-based biometrics has gained attention for its potential to revolutionize user authentication through hands-free interaction, resistance to shoulder surfing, continuous authentication, and revocability. However, current research often relies on single-session or limited-session datasets with fewer than 55 subjects, raising concerns about the generalizability of the findings. To address this gap, we conducted a large-scale study using a public brainwave dataset comprising 345 subjects and over 6,007 sessions (an average of 17 per subject) recorded over five years using three headsets. Our results reveal that deep learning approaches significantly outperform hand-crafted feature extraction methods. We also observe Equal Error Rates (EER) increases over time (e.g., from 6.7% after 1 day to 14.3% after a year). Therefore, it is necessary to reinforce the enrollment set after successful login attempts. Moreover, we demonstrate that fewer brainwave measurement sensors can be used, with an acceptable increase in EER, which is necessary for transitioning from medical-grade to affordable consumer-grade devices. Finally, we compared our results to prior work and existing biometric standards. While our performance is on par with or exceeds previous approaches, it still falls short of industrial benchmarks. Based on the results, we hypothesize that further improvements are possible with larger training sets. To support future research, we have open-sourced our analysis code.
Index Terms:
@@ Abstract
I Introduction
@@
Authentication is the cornerstone of securing digital interactions. While passwords have long been the dominant method, biometrics are increasingly preferred. According to the 2024 FIDO Survey [^1] https://fidoalliance.org/wp-content/uploads/2024/10/Barometer-Report-2024-Oct-29.pdf of 10,000 consumers, 28% preferred biometrics for signing into online accounts, apps, and smart devices, making it the most popular choice. Beyond established methods—such as fingerprint and face recognition—emerging biometric technologies are being explored to enhance usability and security [1, 2]. Brainwave authentication has gained attention for its hands-free operation, which enhances usability in devices like XR systems [3, 4]. It offers resistance to shoulder surfing by relying on distinct neural activity patterns rather than observable actions or physical characteristics [5]. Additionally, brainwave-based systems can enable continuous authentication, improving security by monitoring user identity in active tracking [6]. Furthermore, research indicates that in cases of brainwave sample leakage, changing the stimulus can effectively enable template renewability and revocability, thereby maintaining security [3]. This contrasts with conventional biometric authentication, where compromised raw templates typically cannot be revoked or updated. However, brainwave authentication is still in its early stages, and both permanence and performance remain significant challenges.
@@
Prior studies on brainwave authentication mainly rely on limited public datasets or small, privately collected datasets. Consequently, they either focus on single-session datasets [7, 8, 5, 9] or small-scale multi-session datasets [10, 11, 12, 13]. While single-session studies completely overlook session and temporal effects [14, 15], multi-session studies remain limited to a maximum of 55 subjects [13] and 270 sessions in total [16] (six sessions per subject) so far. These limitations increase the risk of overfitting recognition models and limit the generalizability of results, making it difficult to identify key factors that affect long-term reliability. Importantly, existing multi-session studies rely on a single EEG device, which prevents assessment of hardware variability, although users may access the same profile from different devices in real-world settings. Moreover, the lack of publicly accessible code and data hinders reproducibility and further development, in contrast to the face recognition field, which benefits from open data and collaborative research.
@@
To address this gap and better understand how different parameters affect the performance of brainwave authentication over time, we analyzed the recently published PEERS dataset [17], originally designed for large-scale, longitudinal studies of memory-related electrophysiology. The dataset includes recordings from 345 subjects and 6,007 sessions collected over five years using three different headsets. This dataset contains more than six times the number of subjects and 22 times the number of sessions compared to previous brainwave authentication evaluations. Our contributions are:
  • Benchmark (Sec IV-A): To identify the best pipeline for multi-session brainwave authentication, we conducted a comprehensive evaluation of hand-crafted and deep learning-based feature extraction methods. The results showed that deep learning methods reduced the error rate by up to two times.
@@
  • Analysis (Sec IV-B): To better understand the parameters influencing performance, we investigated the effects of test size, session intervals, and device type. Our analysis revealed several key findings. First, a larger number of subjects in the evaluation set does not necessarily increase authentication error, while smaller datasets lead to higher uncertainty due to limited observations. Second, authentication errors increase over time by 100% after one year. This suggests that updating the enrollment dataset after successful logins may help reduce long-term error rates. Third, visualization of the embedding space shows that samples from the same session and recording device are relatively close. This highlights the influence of sessions and devices on the learning process, likely due to environmental noise, device setup, and physiological or behavioral variations over time. Finally, the analysis of the three devices used for data collection indicates that cross-device authentication is feasible when the feature extractor is trained on subjects who have collected data with both devices.
@@
  • Enhancing Performance (Sec IV-C): We investigated the trade-off between verification time and performance. For example, by increasing the verification time from 1 to 4 seconds, the error rate decreased by approximately 40%. Additionally, we examined the trade-off between enrollment sessions and error rates, observing that after the second enrollment session, the improvement in error rate diminished.
@@
  • Channel Reduction (Sec IV-D): To evaluate the system’s performance under conditions resembling real-world scenarios, we reduced the number of channels in the PEERS dataset—collected using medical-grade devices—from 93 to 14, 7, and 4. These reductions correspond to sensor placements commonly found in consumer-grade EEG devices. The results showed a relatively small increase in error compared to the notable reduction in channels.
  • Industrial Standard Comparison (Sec IV-E): For the first time, we compare brainwave authentication performance against international biometric standards. At required security levels, the false rejection rate (FRR) exceeds acceptable thresholds. However, our results show a logarithmic relationship between error rate and training subjects, suggesting that with more training data, these standards may be met at a practical FRR.
@@
  • Open Sourced Code: We published our source code, which, along with the public dataset, makes results easily reproducible and facilitates future error rate reductions by other researchers  [^2] https://github.com/kit-ps/NeuroShield/.
II Background and Related Work
@@
EEG measures electrical activity produced by neurons firing in the human brain, typically recorded at the scalp using specialized sensors [18, 19]. German physiologist and psychiatrist Hans Berger recorded the first human EEG in 1924 for medical purposes, which remains one of its primary applications [20, 21, 22]. Beyond medical use, EEG technology has advanced to support Brain-Computer Interfaces (BCIs), systems that enable direct communication between the brain and external devices [23]. Recently, EEG data has also been found to include unique features linked to user identity, making it a promising method for user authentication [24]. Brainwave based authentication is hands-free, resistant to shoulder surfing, revocable, renewable, and supports continuous authentication.
@@
Biometric authentication modalities should meet certain requirements to enable widespread usage. These requirements include universality (the modality must be present in all individuals), distinctiveness (it must uniquely identify an individual), permanence (it should remain consistent over time), and collectability (it must be measurable with available technology). Additional factors include performance (accuracy, speed, and resource efficiency), acceptability (user willingness to utilize the system), and circumvention resistance (difficulty in forging or bypassing) [25].
@@
EEG is universal because each person’s brain produces EEG signals [26]. These signals exhibit distinctiveness due to individual brainwave patterns, making them uniquely identifiable [27]. Advances in technology are rapidly improving collectability, with new EEG headsets entering the market annually, offering better signal quality and usability [28]. However, without real-time implementation, it is difficult to comprehensively assess acceptability and circumvention resistance, though some research attempts have been made in this direction [29, 30].
@@
Currently, the primary focus of research is on performance and permanence, as these are challenging aspects of EEG-based authentication. The two are closely interrelated, with performance metrics reflecting the system’s ability to maintain consistent accuracy over time. Factors such as hair growth, slight variations in sensor placement on the scalp, brain states (e.g., stress), and environmental conditions (e.g., noise) can significantly impact both performance and permanence in EEG-based authentication. However, most research studies evaluate the performance of brainwave authentication using single-session datasets (cf. [7, 8, 5, 9]), which neglect session invariance (permanence) completely. This limitation raises concerns that models may overfit to session-specific characteristics rather than capturing subject-unique features. Additionally, some studies (cf. [13, 14]) consider short time intervals, limited to periods of a few days only, which may only partially account for the aforementioned issues. While a limited number of papers explore time intervals spanning more than a week, these studies (cf. [10, 11, 16, 12]) rely on non-public datasets and omit sharing their source code, hindering reproducibility and further research. Moreover, such studies have been conducted with a relatively small sample size and limited sessions (up to 54 subjects [13] and 270 sessions [16]), which increases the risk of overfitting and restricts the generalizability of the results.
@@ III Method
Biometric systems, including brainwave authentication, involve four key stages: data collection, preprocessing, feature extraction, and feature comparison. Like other biometric systems, they operate in two phases: enrollment and verification. In the enrollment phase, brainwave signals are collected from the user to generate a reference profile, known as a template. During verification, performed after enrollment, brainwave signals are captured again and compared with the reference profile to confirm the user’s identity.
TABLE I: Demographic and session distribution of the dataset.
Metric Value
Unique subjects 345
Unique sessions 6007
Average sessions per subject (± SD) 17.41 $\pm$ 10.48
Female 161 (46.7%)
Male 136 (39.4%)
Unspecified sex 48 (13.9%)
Age, years (mean ± SD) 27.0 $\pm$ 15.9
Age, years (range) 17.0–85.0
III-A Dataset Description
@@
Our study utilizes the Penn Electrophysiology of Encoding and Retrieval Study (PEERS) dataset [17], provided by the University of Pennsylvania. Collected between 2010 and 2020, the dataset focuses on exploring EEG correlates of memory processes, particularly during tasks involving word memorization. In these tasks, participants are shown a sequence of words and asked to memorize them as stimulus. Stimulus presentation was synchronized with EEG data collection using event markers. We used the timestamps of word presentations on the screen to segment the EEG data into individual samples. EEG data, recorded as time series of scalp voltage signals, are collected during the sessions at sampling rate of 500 Hz. The dataset includes approximately 8.7 TB of EEG data from 345 subjects, each with more than one session, totaling over 6007 sessions. The participants range in age from 17 to 85 years, with an average age of 26.98 $\pm$ 15.86 years. The dataset includes 161 female and 136 male participants, while gender information for 48 subjects is unavailable (cmp. Table  I). The PEERS dataset is particularly suited for our research as it includes multiple sessions per subject, enabling the investigation of inter-session variability, a critical factor for ensuring the robustness of real-world biometric systems. Data collection was conducted using three distinct medical-grade headsets: the 129-channel Geodesic Sensor Net (GSN 200 model) [^2] https://med.stanford.edu/content/dam/sm/lucasmri/documents/16_0824_EGI_geodesic_sensor_net.pdf , the 129-channel HydroCel Geodesic Sensor Net. [^2] https://med.stanford.edu/content/dam/sm/lucasmri/documents/16_0824_EGI_geodesic_sensor_net.pdf, and the 128-channel BioSemi  [^3] https://www.biosemi.com/headcap.htm headcap using the Biosemi ActiveTwo acquisition system. These diverse devices further enhance the dataset’s relevance by reflecting real-world variability in EEG acquisition systems.
III-B Preprocessing
@@
As the channel layout of the different headsets varied, we unified them by mapping the channels to a well-known extended version [^4] https://mne.tools/stable/auto_tutorials/intro/40_sensor_locations.html of the ten-twenty electrode system [31], which includes 93 channels. Subsequently, the EEG data channels were reordered for each recording to align with the expected order during evaluation across all three devices. Following this, and in line with best practices for preprocessing in brainwave authentication [14, 5, 32, 12], a common average reference (CAR) was applied to reduce spatially correlated noise and improve signal quality. Next, we applied a bandpass filter between 1.0 Hz and 50.0 Hz using a zero-phase finite impulse response (FIR) filter with default parameters, as implemented in MNE-Python [33], to focus on frequency bands that contain EEG data relevant for authentication [34]. As the 50 Hz power-line frequency lies close to the upper end of the preserved frequency band, we additionally applied a notch filter at 50 Hz to suppress potential residual line noise. After filtering, the data was normalized using a robust normalization approach, where the median was subtracted and the result divided by the interquartile range (IQR), mitigating the influence of outliers while preserving the temporal structure of the signal. The data then consisted of 1-second non-overlapping samples aligned to stimulus timestamps. Each sample was stored together with metadata including session date and hardware type.
III-C Feature Extraction
After preprocessing, each sample contains 46,500 values (93 channels × 500 Hz). Due to the high dimensionality and variability of EEG data, direct comparison is not feasible. Therefore, feature extraction is necessary to capture subject-specific information across sessions. Two main approaches exist: the first is handcrafted feature extraction, which uses predefined formulas based on expert knowledge and does not require training data. The second approach uses deep learning techniques, which require a separate training set to learn feature representations. In this method, the feature extractor is trained on labeled data to produce representations that generalize to new subjects.
handcrafted Feature Extraction: We use Power Spectral Density (PSD) to transform EEG signals from the time to frequency domain, capturing patterns linked to brain dynamics. PSD is commonly used in brainwave-based authentication systems [35, 5, 14], applying Fourier transformation to quantify signal power across frequency components [36]. We also extract AutoRegressive (AR) model coefficients, which capture temporal dependencies by modeling each point as a linear combination of previous values. AR features provide compact representations for authentication [5, 37]. Finally, we combine PSD and AR features to incorporate both frequency and temporal information.
Deep learning feature extraction: We applied metric learning, which trains a network to map inputs into a feature space where samples with the same label are close and those with different labels are far [38]. This method uses an encoder to generate embeddings and a loss function that enforces distances based on labels. Metric learning losses are distance-based or softmax-based. Distance-based losses, operating on pairs, triplets, or mined batches, optimize embedding distances directly [39]. Softmax-based losses encourage class separation by clustering embeddings and generally need larger datasets to generalize well [39].